OpenStudy (anonymous):

What is the length of the missing side? A right triangle is shown. One leg is labeled nine. The hypotenuse is labeled fifteen. The second leg is labeled with a question mark.

OpenStudy (anonymous):

Is it a right triangle?

OpenStudy (***[isuru]***):

do u know about Pythagoras theorem ?

OpenStudy (anonymous):

Here are the choices 6 twelve thirty-six square root of three hundred six

OpenStudy (***[isuru]***):



OpenStudy (***[isuru]***):

in a right angle triangle a^2 = b^2 + c^2

OpenStudy (anonymous):

c is usually related to the hypotenuse

OpenStudy (anonymous):

? So would the answer 6\

OpenStudy (***[isuru]***):

in ur problem a = 15 b = 9 u have to find c

OpenStudy (anonymous):



OpenStudy (***[isuru]***):

15^2 = 9^2 + c^2 c^2 = 225 - 81

OpenStudy (***[isuru]***):

solve it for c

OpenStudy (anonymous):

Im so confused the triangkle donyt look like that

OpenStudy (***[isuru]***):

can u post ur triangle... i mean scan it :) ?

OpenStudy (anonymous):

Yes i will try

OpenStudy (anonymous):

\[c^2 = a^2 +b^2\] c = 15 a = b b = ? Plug in the values \[15^2 = 9^2 +b^2\] Square the 15 and 9 \[225 = 81 +b^2\] Subtract 81 from both sides \[144 = b^2\] Take the square root of 144, which is ?

OpenStudy (anonymous):

Sorry a = 9 not b

OpenStudy (anonymous):

rest is correct tho

OpenStudy (anonymous):

|dw:1391720048727:dwl

OpenStudy (anonymous):

opps

OpenStudy (***[isuru]***):

he must keep getting confuse because both of us keep explaining the same thing.... Ryan u may keep the explaining going ...:)

OpenStudy (anonymous):

there is the triangle

OpenStudy (anonymous):

did you see the lunk i sent

OpenStudy (anonymous):

Gave me an error, is it a right triangle ?

OpenStudy (anonymous):

I figured it out thats for the help :)